# Large-Scale Metric Computation in Online Controlled Experiment Platform

### v1

@@ Abstract.
Online controlled experiment (also called A/B test or experiment) is the most important tool for decision-making at a wide range of data-driven companies like Microsoft(Gupta et al., 2018; Kohavi et al., 2013), Google(Tang et al., 2010), Meta(Bakshy et al., 2014), etc. Metric computation is the core procedure for reaching a conclusion during an experiment. With the growth of experiments and metrics in an experiment platform, computing metrics efficiently at scale becomes a non-trivial challenge. This work shows how metric computation in WeChat experiment platform can be done efficiently using bit-sliced index (BSI) arithmetic. This approach has been implemented in a real world system and the performance results are presented, showing that the BSI arithmetic approach is very suitable for large-scale metric computation scenarios.
@@
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
@@ 1. Introduction
  • Experiment data in our situation can be organized into few categories, each category of data is related to WeChat-user activities and attributes. We found that this kind of data follows the Pareto principle(Dunford et al., 2014) (also known as the 80-20 rule) roughly, makes them very efficient when represented by BSI.
@@
Based on these characteristics in our scenarios, this paper first introduce the representation of experiment data, we show that our representation of data using BSI can be very efficient and satisfy the needs of statistical inference. Later we introduce the computation details and conclude that most of the metric computations can be efficiently accomplished by BSI arithmetic. Then we describe our system architecture and present the performance results, shows that how efficient computation and fast interactive query can be achieved in a real world system using reasonable resources.
@@
The rest of the paper is organized as follows. We present the related work in Section 2. Section 3 present the details of the representation of our experiment data by BSI. Algorithms for computing metrics in different scenarios are presented in Section 4. In Section 5, we introduce our system architecture for metric computation. The performance results in Section 6 shows that the efficiency and low latency characteristic of our system. Our conclusions are given in Section 7.
@@ 2.1. Roaring Bitmap
Roaring bitmap was introduced in (Chambi et al., 2016). It is a bitmap scheme using hybrid compression technique that uses both uncompressed bitmaps and packed arrays inside a two-level tree. The roaring bitmap is used to represent a set of 32-bit unsigned integers. A roaring bitmap is a key-value data structure where each key-value pair represents a set $S$ of 32-bit unsigned integers that share the same most significant 16 bits. The key is made of the shared 16 bits, and the value is a container storing the remaining 16 least significant bits for each member of $S$. According to the characteristics of values stored in each container, roaring bitmap uses different container structures for better compress the values. For example, roaring bitmap uses a sorted array when the number of values is less than 4096 in a container (a sparse container). Furthermore, roaring bitmap operations (AND, OR, ANDNOT, XOR, etc.) can be parallelized by exploiting single-instruction-multiple-data (SIMD) instructions(Lemire et al., 2016; Lemire and Boytsov, 2015). It is worth mentioning that the compactness of the integers in a roaring bitmap has great performance impact on bitmap operations due to its design, the denser the bitmap is, the faster the operations are. We use roaring bitmap as a component for building BSI of experiment data.
@@ 2.2. Bit-Sliced Index (BSI)
Bit-sliced sliced index (BSI) was introduced in (O’Neil and Quass, 1997). It is an ordered list of bitmaps, $B^{s},B^{s-1},...,B^{1},B^{0}$, and is used to represent the values (normally non-negative integers, and we use BSI to represent non-negative numeric values in our system) of some column $C$ of a table $T$ (although the column $C$ might be calculated values associated with rows of $T$, and have no physical existence in the table). The bitmaps $B^{i},0\leq i\leq s$ are called bit-slices, and their bit-values are defined this way:
@@ 2.3. BSI Arithmetic
It is easy to see that the computational complexity of addition, subtraction, comparsion operators(<,=,<,=$\neq$< , = , ≠), in-BSI aggregate functions and range searches grows linearly as the number of non-empty bit slices increases. While the complexity of general multiplication of BSI is $O(s_{1}s_{2})$, $s_{1}$ and $s_{2}$ are the numbers of non-empty bit slices in each BSI, which seems to be slower. Fortunately, in our scenarios we only need the multiplication with one of the operators being binary, which makes the complexity of multiplication also linear.
@@ 3.3. Bucketing and Statistical Inference
We suggest that it is reasonable to assume the randomization units satisfy the stable unit treatment value assumption (SUTVA)(Rubin, 1980, 1986), which requires that ”the observation of potential outcome on one unit should not be affected by the particular assignment of treatments to the other units”. It is an assumption about independence among randomization units, but goes beyond the concept of independence. If the assumption is not satisfied, for example, in a user-randomized experiment, a user who is exposed to experiment strategy B will have different stay-time, depends on the strategy exposure of other users, then the stay-time computed by users in strategy B with partial users in it can not stand for the real effect of strategy B when it is full launched, which makes the experiment conclusion invalid.
Based on the ”independence” property of the randomization units, we introduce a bucketing procedure for estimating variance and covariance of metrics efficiently. Similar to what we did for segmentation, we introduce a deterministic randomization process for assigning all randomization units into different buckets 0, 1, …, 1023, and we compute the metric value for each bucket separately. It can be seen as generating 1024 independent replicates of the metric values of an experiment strategy, and then bootstrapping(Johnson, 2001) the variance and covariance of metrics. We proved that the bucket based method of estimating the variance and covariance of metrics is correct in theory, and it makes the statistical inference procedures more efficient, check (Xiong et al., 2021) for more details.
It is worth mentioning that the randomization unit and the analysis unit are the same in most of the experiments, in which case the segmentation and bucketing become the same procedure.
@@ 3.5. Efficiency of BSI Representation
An important property of BSI with roaring bitmaps is that the data will be well compressed when the most of the values of BSI are concentrated in relatively small range and the position encoding is compact, because the roaring bitmap will compress the zeros of binary representation of each value and keep bitmap operations available on compressed data at the same time. This kind of compression is different from the data storage compression, it compresses the data size processed by CPU when executing the related operations on the data, for example, for raw data with two columns of size 1GB we need process 1GB data in CPU when adding these two columns, while for compressed BSI data we maybe only need process 100MB data in CPU, which improves the performance significantly if the addition operation on BSI are efficient enough compared to the addition operation on raw data.
We evaluate the performance of BSI representation in section 6.
@@ 4.1.2. Filter
Part of the values in BSI can be filtered out with predicates on BSIs, for example, we can select the expose information of a strategy whose analysis units are first exposed between the 2nd day and 5th day (the symbol * is for multipliation of BSI, and the comparison is also performed on BSI):
@@ 4.3. Pre-Experiment Computation
In (Deng et al., 2013) a technique was introduced that uses pre-experiment data to reduce the variance in experimentation metrics. We also implemented the method in our system. The computation of pre-experiment is similar to the scorecard computation, except that the expose-log is joined with the metric-log previous to the experiment start date. We also demonstrated the single-day pre-experiment computation for one strategy-metric pair just like done before.
@@
        ’expt_start_date-C’ AND ’expt_start_date-1’
@@
It is worth noting that $C$ successive days of metric-log have been aggregated by $sumBSI$ first, this procedure can be accelerated by pre-aggregating the metric-log with a tree data structure. As illustrated in Figure 6, each non-leaf node of the tree is merged from its two children by some aggregate function over BSIs. We can compute aggregation of successive days more efficiently by merging less nodes in the tree, for example, performing $sumBSI$ for day 1 to day 7 can be accomplished by merging 3 nodes (1234, 56, 7) in the tree instead of 7 nodes.
@@ 5.1. Overall Architecture
The architecture of the metric computation is illustrated in Figure 7 in the next page. The raw expose log and the raw metric/dimension log are first processed, converted to BSI representations and stored on a distributed data warehouse system. Routine computations like the scorecard computation and the pre-experiment computation are executed by pre-compute pipeline using Apache Spark(Zaharia et al., 2010, 2012). Ad-hoc queries like deep dive analysis are processed by Clickhouse (an open-source columnar database management system for online analytical processing).
@@ 6.1. BSI Representation Evaluation
In subsection 3.5 We use 5890 real world user-level metrics in our system for analyzing the characteristics of the data, and conclude that we can compress the data through the BSI representation and keep the operations on it available without decompression.
In this section we evaluation the storage efficiency of the BSI representation. We choose 105 user-level core metrics, these core metrics are the most important metrics of one of the vital business scenarios in WeChat. The value range cardinalities distribution of the these metrics in one day are illustarted in Table 3, the table shows that these metrics tend to have bigger value ranges compared to the 5890 metrics, and they should be enough for evaluation usage (the performance will be better for the 5890 metrics). We do not use more metrics here, because too many metrics will lead to too much storage and computational cost.
@@
Metric Rows Original Size Value Range
A 316 million 140 MB (0, 1]
B 34 million 86 MB (0, 50]
C 510 million 2 GB (0, 21600]
Next we evaluate the computational cost of the BSI representation. We choose 3 typical metrics with different value ranges and rows. The details of the 3 metrics are illustrated in Table 5, metric C has a big value range and the maximum number of rows, thus has a maximum size of 2GB in one day, metric B has a bigger value range than metric A, but also has much less rows than metric A, and the final size of metric B is smaller than metric A. For each metric, we use one CPU core for calculating the sum of metric values for each user in two days (that is performing $sumBSI$ for two BSI or aggregating by user-id for normal format in each segment). The evaluation program is single threaded and written in C++, compiled with $-O3$ optimization option, and runs repeatedly on a machine with a 2.4GHZ CPU. The performance results are listed in Table 6, it shows that the computation of BSI representation is much more efficient than that with the normal format. We conclude that the performance gain of the BSI computation mainly comes from the compression of data when processed in CPU and the SIMD instructions when performing BSI operations.
@@
Metric Normal Format BSI Format
A 59.2 seconds 0.6 seconds
B 7.3 seconds 1.3 seconds
C 94.3 seconds 10.5 seconds
@@ 6.2. Pre-Computation Evaluation
Pre-computation performance is also evaluated on the 105 core metrics. There are about 240,000 strategy-metric pairs to be computed in November 21, 2023 in our system, these pairs involves about 8,500 strategies and each strategy contains an average of 21 million exposed users. We evaluate the CPU hours consumed for computing the scorecard results of these pairs using two methods. The BSI based method is described in subsection 4.2, while the normal representation based method is described as follow (this is also the method used in our system before the deployment of the BSI method):
@@
The results are illustrated in Table 7, the CPU hours of the BSI based method are about a quarter of that with the normal format.

### v2

@@ Abstract.
Online controlled experiment (also called A/B test or experiment) is the most important tool for decision-making at a wide range of data-driven companies like Microsoft, Google, Meta, etc. Metric computation is the core procedure for reaching a conclusion during an experiment. With the growth of experiments and metrics in an experiment platform, computing metrics efficiently at scale becomes a non-trivial challenge. This work shows how metric computation in WeChat experiment platform can be done efficiently using bit-sliced index (BSI) arithmetic. This approach has been implemented in a real world system and the performance results are presented, showing that the BSI arithmetic approach is very suitable for large-scale metric computation scenarios.
@@
PVLDB, 17(12): 4014 - 4024, 2024.
doi:10.14778/3685800.3685823 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 12 ISSN 2150-8097.
doi:10.14778/3685800.3685823
@@ 1. Introduction
Online controlled experiment (A/B tests) plays a central role in data-driver decision-making in the Internet industry (Gupta et al., 2018; Kohavi et al., 2013; Tang et al., 2010; Bakshy et al., 2014). With the collection of massive data from users during the experiments, efficient data processing and analyzing is crucial for reaching conclusions of experiments, and metric computation is the most important part of it.
@@
  • Experiment data in our situation can be organized into few categories, each category of data is related to WeChat-user activities and attributes. We found that this kind of data follows the Pareto principle (Dunford et al., 2014; Newman, 2005) (also known as the 80-20 rule) roughly, makes them very efficient when represented by BSI.
@@
Based on these characteristics in our scenarios, this paper first introduce the representation of experiment data, we show that our representation of data using BSI can be very efficient and satisfy the needs of statistical inference. Later we introduce the computation details and conclude that most of the metric computations can be efficiently accomplished by BSI arithmetic. Then we describe our system architecture and present the performance results, shows that how efficient computation and fast interactive query can be achieved in a real world system using reasonable resources. At last, we discuss the expressive power and limitation of the proposed method.
@@
The rest of the paper is organized as follows. We present the related work in section 2. section 3 present the details of the representation of our experiment data by BSI. Algorithms for computing metrics in different scenarios are presented in section 4. In section 5, we introduce our system architecture for metric computation. The performance results in section 6 shows that the efficiency and low latency characteristic of our system. The expressive power and limitation are discussed in section 7. Our conclusions are given in section 8.
@@ 2.1. Roaring Bitmap
Roaring bitmap was introduced in (Chambi et al., 2016; Lemire et al., 2016b). It is a bitmap scheme using hybrid compression technique that uses both uncompressed bitmaps and packed arrays inside a two-level tree. The roaring bitmap is used to represent a set of 32-bit unsigned integers. A roaring bitmap is a key-value data structure where each key-value pair represents a set $S$ of 32-bit unsigned integers that share the same most significant 16 bits. The key is made of the shared 16 bits, and the value is a container storing the remaining 16 least significant bits for each member of $S$. According to the characteristics of values stored in each container, roaring bitmap uses different container structures for better compress the values. For example, roaring bitmap uses a sorted array when the number of values does not exceed 4096 in a container (a sparse container). Furthermore, roaring bitmap operations (AND, OR, ANDNOT, XOR, etc.) can be parallelized by exploiting single-instruction-multiple-data (SIMD) instructions (Lemire et al., 2016a; Lemire and Boytsov, 2015). It is worth mentioning that the compactness of the integers in a roaring bitmap has great performance impact on bitmap operations due to its design, the denser the bitmap is, the faster the operations are. We use roaring bitmap as a component for building BSIs of experiment data.
@@ 2.2. Bit-Sliced Index (BSI)
Bit-sliced index (BSI) was introduced in (O’Neil and Quass, 1997). It is an ordered list of bitmaps, $B^{s},B^{s-1},...,B^{1},B^{0}$, and is used to represent the values (normally non-negative integers, and we use BSI to represent non-negative numeric values in our system) of some column $C$ of a table $T$ (although the column $C$ might be calculated values associated with rows of $T$, and have no physical existence in the table). The bitmaps $B^{i},0\leq i\leq s$ are called bit-slices, and their bit-values are defined this way:
@@ 2.3. BSI Arithmetic
It is easy to see that the computational complexity of addition, subtraction, comparison operators(<,=,<,=$\neq$< , = , ≠), in-BSI aggregate functions and range searches grows linearly as the number of non-empty bit slices increases. While the complexity of general multiplication and division of BSI are $O(s_{1}s_{2})$, $s_{1}$ and $s_{2}$ are the numbers of non-empty bit slices in each BSI, which seems to be slower. Fortunately, in the scenarios of this paper we only need the multiplication with one of the operators being binary, which makes the complexity also linear.
@@ 3.3. Bucketing and Statistical Inference
We suggest that it is reasonable to assume the randomization units satisfy the stable unit treatment value assumption (SUTVA) (Rubin, 1980, 1986; Sekhon, 2008), which requires that ”the observation of potential outcome on one unit should not be affected by the particular assignment of treatments to the other units”. It is an assumption about independence among randomization units, but goes beyond the concept of independence. If the assumption is not satisfied, for example, in a user-randomized experiment, a user who is exposed to experiment strategy $B$ will have different stay-time, depends on the strategy exposure of other users, then the stay-time computed by users in strategy $B$ with partial users in it can not stand for the real effect of strategy B when it is full launched, which makes the experiment conclusion invalid.
Based on the ”independence” property of the randomization units, we introduce a bucketing procedure for estimating variance and covariance of metrics efficiently. Similar to what we did for segmentation, we introduce a deterministic randomization process for assigning all randomization units into different buckets 0, 1, …, 1023, and we compute the metric value for each bucket separately. It can be seen as generating 1024 independent replicates of the metric values of an experiment strategy, and then bootstrapping (Johnson, 2001; Mooney et al., 1993) the variance and covariance of metrics. We proved that the bucket-based method of estimating the variance and covariance of metrics is correct in theory, and it makes the statistical inference procedures more efficient, check (Xiong et al., 2021) for more details.
It is worth mentioning that the randomization unit and the analysis unit are the same in most of the experiments, in which case the segmentation and bucketing can be the same procedure.
@@ 3.5. Efficiency of BSI Representation
An important property of BSI with roaring bitmaps is that the data will be well compressed when the most of the values of BSI are concentrated in relatively small range and the position encoding is compact, because the roaring bitmap will compress the zeros of binary representation of each value and keep bitmap operations available on compressed data at the same time. This kind of compression is different from the data storage compression, it compresses the data size processed by CPU when executing the related operations on the data, for example, for raw data with two columns of size 1GB we need process 1GB data by CPU when adding these two columns, while for compressed BSI data we maybe only need process 100MB data by CPU, which improves the performance significantly if the addition operation on BSI are efficient enough compared to the addition operation on raw data.
We evaluate the performance of BSI representation in section 6 with more details.
@@ 4.1.2. Filter
Part of the values in BSI can be filtered out with predicates on BSIs, for example, we can select the expose information of a strategy whose analysis units are first exposed between the 2nd day and 5th day (the symbol * is for multiplication of BSI, and the comparison is also performed on BSI):
@@ 4.3. Pre-Experiment Computation
In (Deng et al., 2013) a technique was introduced that uses pre-experiment data to reduce the variance in experimentation metrics. We also implemented the method in our system. The computation of pre-experiment is similar to the scorecard computation, except that the expose-log is joined with $C$ successive days of the metric-log previous to the experiment start date. We also demonstrated the single-day pre-experiment computation for one strategy-metric pair just like done before.
@@
        (expt_start_date - C) AND (expt_start_date - 1)
@@
It is worth noting that $C$ successive days of metric-log are aggregated by $sumBSI$ first, this procedure can be accelerated by pre-aggregating the metric-log with a tree data structure. As illustrated in Figure 6, each non-leaf node of the tree is merged from its two children by some aggregate function over BSIs. We can compute aggregation of successive days more efficiently by merging less nodes in the tree, for example, performing $sumBSI$ for day 1 to day 7 can be accomplished by merging 3 nodes (1234, 56, 7) in the tree instead of 7 nodes.
@@ 5.1. Overall Architecture
The architecture of the metric computation is illustrated in Figure 7 in the next page. The raw expose log and the raw metric/dimension log are first processed, then converted to BSI representations and stored on a distributed data warehouse system. Routine computations like the scorecard computation and the pre-experiment computation are executed by pre-compute pipeline using Apache Spark (Zaharia et al., 2010, 2012). Ad-hoc queries like deep dive analysis are processed by Clickhouse (Imasheva et al., 2020) (an open-source columnar database management system for online analytical processing).
@@ 6.1. BSI Representation Evaluation
In subsection 3.5 We use 5890 real world user-level metrics in our system for analyzing the characteristics of the data, and conclude that we can compress the data through the BSI representation and keep the operations on it available without decompression. In this subsection we evaluation the storage and computational efficiency of the BSI representation, or BSI format.
6.1.1. Storage Evaluation
We choose 105 user-level core metrics, these core metrics are the most important metrics of one of the vital business scenarios in WeChat. The value range cardinalities distribution of the these metrics in one day are illustrated in Table 3, the table shows that these metrics tend to have bigger value ranges compared to the 5890 metrics, and they should be enough for evaluation usage (the performance will be better if the evaluation is carried out with the 5890 metrics). We do not use more metrics here, because too many metrics will lead to too much storage and computational cost.
@@ 6.1.1. Storage Evaluation
Metric Rows Normal Size BSI Size Value Range
A 316 million 2.9 GB 140 MB (0, 1]
B 34 million 324 MB 86 MB (0, 50]
C 510 million 4.7 GB 2.0 GB (0, 21600]
6.1.2. Computational Evaluation
Next we evaluate the computational cost of the BSI representation. We choose 3 typical metrics with different value ranges and rows. Raw data of each metric in one day is represented as rows of (segment-id, user-id, metric-value) with normal format, and there is at most one row for each user in one day (that is, aggregated by user-id within one day). Raw data of the metrics is converted to BSI representation for evaluation. The details of the 3 metrics in one day are illustrated in Table 5, The size of data that processed by CPU with normal format (Normal Size column) is only affected by number of rows, while the size of BSI is determined by number of rows and value range (value distribution). Metric $C$ has a big value range and the maximum number of rows, thus has a maximum BSI size in one day, metric $B$ has a bigger value range than metric $A$, but also has much less rows than metric $A$, and the final BSI size of metric $B$ is smaller than metric $A$.
For each metric, we use one CPU core for calculating the sum of metric values for each user in two days (that is, within each segment, performing $sumBSI$ for two BSIs or aggregating two days raw data by user-id). The evaluation program is single threaded and written in C++, compiled with $-O3$ optimization option, and runs 10 times repeatedly on a machine with a 2.4GHZ CPU. The performance results are listed in Table 6, it shows that the computation of BSI representation is much more efficient than that with the normal format, the main reason is that the size of data processed by CPU with BSI representation is much more smaller than that with normal format, and the SIMD instructions also play an important role in performance improvement.
We conclude that the performance gain of the BSI computation mainly comes from the compression of data when processed by CPU and the SIMD instructions when performing BSI operations.
@@ 6.1.2. Computational Evaluation
Format Metric A Metric B Metric C
Normal 59.2 seconds 7.3 seconds 94.3 seconds
BSI (no SIMD) 1.2 seconds 3.8 seconds 18.3 seconds
BSI (AVX2) 0.6 seconds 1.3 seconds 10.5 seconds
6.1.3. Extra Cost in Log Processing
We evaluated the extra computational cost in log processing derived from conversion to the BSI format in the same settings. For each metric, we converted a day’s worth of data from the normal format to the BSI format using a straightforward method, extracting bits from each metric value and setting these bits into corresponding bitmaps. We also observed that the conversion process can be more efficient when the data is pre-sorted by user-id in the normal format. By splitting the pre-sorted rows into blocks and processing each block with neighbouring user-ids, we set bits extracted from metric values in the block into adjacent containers in the roaring bitmap. This approach achieves better cache locality, enhancing the overall efficiency of the conversion. The results are illustrated in Table 7. It is easy to see that the conversion cost is mainly related to the number of rows and the value range of the metric, more rows and bigger value range cost more CPU time. In our situation, most metrics have relatively small value ranges, and the total computation cost (conversion and BSI computation) is still much lower than that with normal format in our scenarios.
Furthermore, we have monitored the overall cost (CPU hours and latency) of log processing in our real system, and the results show that the log processing cost remains almost the same as before, the reason is that there are more IO-bound operations in log processing, so the BSI conversion cost does not become a bottleneck in this procedure.
Figure 9. Convert back to Normal Format within a Container
6.1.4. Cost of Converting Back to the Normal Format
We also evaluated the cost of converting data back to the normal format in the settings above. For each metric, we converted a day’s worth of data from the BSI format to the normal format using both a straightforward method and a per-bitmap method. Using the straightforward method, we collect the bits of the metric-value for a user from each bitmap and combine them into a single value. The problem of the straightforward method is that we need to check all the bitmaps for each metric value, even if there is only one bit in the value. Additionally, the scattered bitmap access reduces the cache hit rate. To resolve the problem, we extract the bits in a per-bitmap manner, as illustrated in Figure 9, for each bitmap, bits identified by a mask (indicating the user-ids of interest) within a container are first extracted and then set into corresponding positions of metric values, the output is then generated by retrieving the metric values corresponding to the mask, this process is repeated for each container of the roaring bitmap. This approach avoids the extra cost of processing zero bits in metric values and achieves better cache locality. Additionally, the bit extraction process can also be accelerated by SIMD instructions when necessary. The evaluation results, illustrated in Table 8, shows that the conversion back process is very efficient with the per-bitmap method. Furthermore, the output of the conversion is naturally sorted by user-id within the mask, and can be efficiently converted to the BSI format again.
Table 7. Average Time of Converting to BSI format
Method Metric A Metric B Metric C
Straightforward 4.2 seconds 0.9 seconds 34.9 seconds
Pre-sorted 3.9 seconds 0.8 seconds 12.9 seconds
Table 8. Average Time of Converting back to Normal Format
Method Metric A Metric B Metric C
Straightforward 44.6 seconds 10.0 seconds 164.6 seconds
Per-bitmap 2.0 seconds 1.1 seconds 8.7 seconds
@@ 6.2. Pre-Computation Evaluation
Pre-computation performance is also evaluated on the 105 core metrics. There are about 240,000 strategy-metric pairs to be computed in November 21, 2023 in our system, these pairs involves about 8,500 strategies and each strategy contains an average of 21 million exposed users. We evaluate the CPU hours consumed for computing the scorecard results of these pairs using two methods. The BSI-based method is described in subsection 4.2, while the normal representation based method is described as follow (this is also the method used in our system before the deployment of the BSI method):
@@
The results are illustrated in Table 9, the CPU hours of the BSI-based method are about a quarter of that with the normal format.